When you add 12 to a number and then divide the sum by 13, you get the same result as when you subtract 13 from the number and then divide the difference by 12. What's the number?
Answer by stanbon(75887)

When you add 12 to a number and then divide the sum by 13, you get the same result as when you subtract 13 from the number and then divide the difference by 12. What's the number?
----
Let the number be "x":
Equation:
(x+12)/13 = (x-13)/12
-----
Cross-multiply:
12x + 144 = 13x - 169
---
x = 313
===================
Cheers,
Stan H.
============